OpenStudy (anonymous):

solve r=1/2m^2p for p

OpenStudy (solomonzelman):

Example: \(\color{#000000 }{ \displaystyle t=\frac{2}{5}k^3z }\) solve for z. \(\color{#000000 }{ \displaystyle t\color{red}{\times \frac{5}{2}}=\frac{2}{5}k^3z\color{red}{\times \frac{5}{2}} }\) \(\color{#000000 }{ \displaystyle \frac{5}{2}t=k^3z }\) \(\color{#000000 }{ \displaystyle \frac{5}{2}t\color{red}{\div k^3}=k^3z\color{red}{\div k^3} }\) \(\color{#000000 }{ \displaystyle \frac{5}{2}t/k^3=z }\)

OpenStudy (anonymous):

i still dont get it :(